OpenStudy (anonymous):

Solve: 2x + 36 + 40 = 180

OpenStudy (anonymous):

2x = 180 - 36 - 40

OpenStudy (anonymous):

2x = 104

OpenStudy (anonymous):

x = 104/2 Therefore , x = 52